ANALYSIS OF THE ROTATIONAL SPECTRA OF 2,3,4,5,6-PENTAFLUOROTOLUENE AND 1-CHLORO-2,3,4,5,6-PENTAFLUOROBENZENE

\underline{ASHLEY A. OSTHOFF}, REBECCA A. PEEBLES, SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920; GARRY S. GRUBBS II, STEPHEN A. COOKE, Department of Chemistry, University of North Texas, PO Box 305070, Denton, TX 76203; BROOKS H. PATE, JUSTIN L. NEILL, MATT T. MUCKLE, Department of Chemistry, University of Virginia, McCormick Rd., PO Box 400319, Charlottesville, VA 22904.

The microwave spectra of two substituted pentafluorobenzenes have been obtained. 2,3,4,5,6-Pentafluorotoluene was measured using the FTMW spectrometer at Eastern Illinois University and the chirped-pulse FTMW spectrometer at University of North Texas. The heavy atom structure has been obtained from the assigned ^{13}C transitions and is in reasonable agreement with ab initio calculations at the MP2/6-311++G(2d, 2p) level. The ground state rotational constants are A = 1036.61253(10) MHz, B = 1030.94126(10) MHz, and C = 516.92062(9) MHz, and the single dipole moment component is μ_{b} = 1.98(17) D. Very small splittings for many of the assigned transitions and multiple, as yet unassigned, lines were presumably due to excited torsional states of the methyl group. In a related study, the microwave spectrum of 1-chloro-2,3,4,5,6-pentafluorobenzene was obtained for both the ^{35}Cl and the ^{37}Cl isotopologues using the chirped-pulse microwave spectrometer at University of Virginia. The preliminary ground state rotational constants for this compound are A = 1028.5403(14) MHz, B = 751.8198(3) MHz and C = 434.3533(4) MHz for ^{35}Cl and A = 1028.5435(7) MHz, B = 734.4786(2) MHz and C = 428.5082(2) MHz for ^{37}Cl. Initial fits of the nuclear quadrupole coupling constants give χ_{aa} = −79.512(15) MHz, χ_{bb} = 43.593(8) MHz, χ_{cc} = 35.92(2) MHz for the ^{35}Cl species and χ_{aa} = −62.68(2) MHz, χ_{bb} = 34.38(4) MHz, χ_{cc} = 28.29(17) MHz for the ^{37}Cl species. These results will be compared with pentafluorotoluene to observe the effects on the structure of the benzene ring when substituting a chlorine atom for a methyl group.